# Nonreciprocal quantum phase transition in a spinning microwave magnonic system

### v1

@@ Abstract
We propose how to achieve nonreciprocal quantum phase transition in a spinning microwave magnonic system composed of a spinning microwave resonator coupled with an yttrium iron garnet sphere with magnon Kerr effect. Sagnac-Fizeau shift caused by the spinning of the resonator brings about a significant modification in the critical driving strengths for second- and one-order quantum phase transitions, which means that the highly controllable quantum phase can be realized by the spinning speed of the resonator. More importantly, based on the difference in the detunings of the counterclockwise and clockwise modes induced by spinning direction of the resonator, the phase transition in this system is nonreciprocal, that is, the quantum phase transition occurs when the system is driven in one direction but not the other. Our work offers an alternative path to engineer and design nonreciprocal magnonic devices.
@@ I Introduction
Quantum phase transition (QPT) has attracted considerable attention for its crucial role in quantum physics and potential applications in modern quantum technologies 1 ; 1a ; 2 ; 2a ; 2b ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 10 ; 11 ; 12 ; 13 ; 14 . Different from the classical phase transition happening in a finite temperature, this type of phase transition is mainly originated from the quantum fluctuation at the critical parameter regime and can take place between two stable phases accompanied by spontaneous symmetry breaking as the temperature tends to the absolute zero of temperature. Usually, a QPT occurs in the thermodynamic limit involving a infinite number of system components. Although recent advances about QPTs have indicated that the thermodynamic limit may be no longer a requirement 15 ; 16 ; 17 , the QPT observed in infinite-component system is still one of the most fascinating topics. Recently, it is interesting that that a squeezed light driving high-quality cavity can give rise to QPT in a dissipative qubit collective system without the ultra-strongly nonlinear coupling between the atoms and the field 18 . Additionally, the QPTs have also been studied in dissipative systems with other types of nonlinearity, such as the Kerr nonlinearity 19 ; 20 . Among them, it was confirmed that a QPT can emerge in a Kerr-nonlinear oscillator subject to two-photon driving and single-photon loss. Moreover, based on the magnon Kerr effect owing to the magnetocrystalline anisotropy in the ferromagnetic yttrium-iron-garnet sphere, the studies on QPT have made progress in hybrid cavity magnonic systems in recent years. In Ref. 21 , the authors studied the parity-symmetry-breaking QPT in a single-cavity magnonic system driven by a parametric field. This system can show a rich phase diagram, including the parity-symmetric phase (PSP), parity-symmetry-broken phase (PSBP), and bistable phase (BP). Subsequently, a controllable QPT was also proposed in a double-cavity magnonic system 22 . In particular, they demonstrated that the phase transition in one cavity can be precisely and efficiently controlled by adjusting the parameters of the other cavity. Nonetheless, whether the controllability of QPT can be realized only in a single-cavity magnonic system is still a question worth exploring.
@@
Nonreciprocal physics, referring to the phenomenon that a system displays different responses in opposite directions, has witnessed rapid advances in recent years. Up to now, the study of nonreciprocity has been extended to many scientific branches, such as optics 23 ; 24 , acoustics 25 ; 26 ; 27 , thermodynamics 28 . As an important application, nonreciprocal devices allow signals to propagate in one direction and prevents signals from propagating in the opposite direction, and therefore plays a key role in backaction-immune communications. Traditional nonreciprocal devices break Lorentz reciprocity to achieve nonreciprocity mainly by employing the Faraday effect in magneto-optical crystal materials 29 ; 30 , but such devices needs the magnetic materials that are bulky, costly, and unsuitable for on-chip integration due to the high susceptibility to external magnetic field interference. To circumvent these obstacles, a variety of nonreciprocity schemes have been proposed based on nonlinear optics 31 ; 32 ; 32a ; 32b , optomechanics 33 ; 34 ; 35 , non-Hermitian optics 36 ; 37 ; 38 ; 39 . Especially, a recent experiment confirmed that an optical diode with 99.6% isolation has been realized by using a spinning resonator 40 , in which the nonreciprocity is caused by the Fizeau shift of circulating lights. Subsequently, there is increasing interests in the connection of the nonreciprocity with other quantum effects, such as nonreciprocal quantum blockades 41 ; 42 ; 42a ; 43 ; 44 ; 45 ; 46 ; 47 ; Wang22 ; 48 , quantum entanglement 49 ; 50 ; 51 ; 52 , mechanical squeezing 53 ; 54 ; 55 , phonon and magnon laser 56 ; 57 ; 58 , and sideband responses 59 ; 60 . Very recently, one particular concern is that nonreciprocal superradiant phase transitions have been reported in a cavity QED system 61 .
Here we propose how to achieve an nonreciprocal QPT in a single spinning microwave resonator coupled to an yttrium iron garnet (YIG) sphere with magnon Kerr nonlinearity. We show that, compared with the case of double-cavity, the QPT can be efficiently controlled only by adjusting the angular speed of the resonator without introducing an auxiliary resonator. Physically, the value of the order parameter for characterizing QPT is strongly dependent on the detuning between the cavity field and the driving laser. Meanwhile, the Sagnac effect caused by the spinning of the resonator can result in different influences on the detuning of the counterclockwise (CCW) and clockwise (CW) fields compared to the driving laser. Therefore, nonreciprocal QPT occurs by reason of the significant difference in the mean magnon number for driving the resonator from the opposite directions. Specifically, our system can display a rich phase diagram containing PSP, PSBP, and BP. Furthermore, we show that the system can experience a second-order phase transition from a PSP to a PSBP or a first-order phase transition from a PSP to a BP when increasing the driving strength beyond a critical value under different initial conditions. More importantly, in comparison to the stationary case, the tunable nonreciprocal QPT can be achieved by the Sagnac effect. Furthermore, we investigate the different behaviors of the mean magnon number and correlation fluctuation in the vicinity of the critical point and find that these behaviors can be used to identify different phase transitions. In addition, an isolation parameter is introduced to quantitatively describe the nonreciprocal degree. The results indicate again that the QPT can be readily switched on and off by adjusting the driving direction or the spinning direction of the resonator.
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the theoretical model and give the dynamics and the steady-state solutions of the system through Heisenberg-Langevin approach. In Sec. III, the impact of the Sagnac-Fizeau shift and the rotating direction of the resonator on the nonreciprocal behaviors of the QPT are studied in detail. Finally, we summarize our results in Sec. IV.
@@ II The model
Figure 1: (a) Schematic of the system. A YIG sphere with magnon Kerr nonlinearity is placed inside a spinning microwave resonator driven by a squeezed light generated by a flux-driven JPA and simultaneously biased by a uniform magnetic field along the z direction. (b) Frequencies of the microwave resonator. When the microwave resonator rotates in the fixed counterclockwise (CCW) direction, the resonance frequencies of the CCW mode and clockwise (CW) mode will experience different Fizeau shifts, where the frequencies $\omega_{a}+\left|\Delta_{F}\right|$ and $\omega_{a}-\left|\Delta_{F}\right|$ denote for the CCW and CW modes, respectively.
As schematically shown in Fig. 1, we consider a cavity-magnon system which is composed of a spinning microwave resonator coupled to a YIG sphere via the magnetic dipole interaction. While the resonator is driven by a weak squeezed vacuum field generated by a flux-driven Josephson parametric amplifier (JPA) and the magnetocrystalline anisotropy results in the Kerr nonlinear interaction among magnons in the YIG sphere. The Hamiltonian of the total system reads ($\hbar=1$)
@@
where $a$ $\left(a^{\dagger}\right)$ and $m$ $\left(m^{\dagger}\right)$ are, respectively, the annihilation (creation) operators of the cavity and magnon modes with the corresponding resonance frequencies $\omega_{a}$ and $\omega_{m}$. The microwave resonator is parametrically driven with driving amplitude $G$ and pump frequency $\omega_{d}$. $K$ denotes the nonlinear coefficient of the magnon Kerr effect and $J$ is the strength of the cavity-magnon interaction. For a resonator spinning at an angular velocity $\Omega$, the frequencies of the CCW and CW modes experience Sagnac-Fizeau shift, i.e., $\omega_{a}\rightarrow\omega_{a}+\Delta_{F}$, with
@@
in which $n$ is the refractive index, $r$ is the radius of the resonator, and $\lambda$ $(c)$ is the wavelength (speed) of the light in vacuum. The dispersion term $dn/d\lambda$, characterizing the relativistic origin of the Sagnac effect, is relatively small ($\sim$1%) and thus can be ignored. In the rotating frame with respect to the driven frequency $\omega_{d}$, the Hamiltonian of the system becomes
$H$ $=$ $\tilde{\Delta}_{a}a^{\dagger}a+\Delta_{m}m^{\dagger}m+\frac{K}{2}m^{\dagger}m^{\dagger}mm$ (3)
@@
with $\tilde{\Delta}_{a}=\Delta_{a}+\Delta_{F},\Delta_{a}=\omega_{a}-\omega_{d}/2,\Delta_{m}=\omega_{m}-\omega_{d}/2$.
By considering dissipation and noise effects in the Hamiltonian (3), the quantum Langevin equations (QLEs) describing the dynamics of the hybrid system can be written as
$\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)a-iJm-iGa^{\dagger}+\sqrt{2\kappa}a_{in},$
$\dot{m}$ $=$ $-i\left(\Delta_{m}-i\gamma\right)m-iKm^{\dagger}mm-iJa+\sqrt{2\gamma}m_{in},$ (4)
where $\kappa$ $\left(\gamma\right)$ and $a_{in}$ $\left(m_{in}\right)$ are the damping rate and the zero-mean input noise operator for the cavity (magnon) mode, respectively. Furthermore, a linearization procedure can be used to simplify the physical model. We expand the operators $a$ and $m$ in Eq. (4) as a sum of their expectation values and small quantum fluctuations, $a=A+\delta a$ $\left(m=M+\delta m\right)$. By using the mean-field approximation, it follows from Eq. (4) that the dynamical equations for the expectation values are
@@
In the steady-state case $\left(\dot{A}=\dot{M}=0\right)$, solving Eq. (5) gets three solutions for the mean magnon number $\left|M\right|^{2}$, i.e.,
$\left|M\right|_{0}^{2}=0,\text{ }\left|M\right|_{\pm}^{2}=\frac{-\tilde{\Delta}_{m}\pm\sqrt{\xi^{2}G^{2}-\tilde{\gamma}^{2}}}{K},$ (6)
with $\tilde{\Delta}_{m}=\Delta_{m}-\beta\tilde{\Delta}_{a}$, $\tilde{\gamma}=\gamma+\beta\kappa$, $\beta=J^{2}/\left(\tilde{\Delta}_{a}^{2}+\kappa^{2}-G^{2}\right)$. According to the first equation in Eq. (5), the steady-state photon occupation can be expressed as
$\left|A\right|_{0}^{2}=0,\text{ }\left|A\right|_{\pm}^{2}=\frac{\left(\tilde{\Delta}_{m}+\left|M\right|_{\pm}^{2}\right)^{2}+\gamma^{2}}{g^{2}}\left|M\right|_{\pm}^{2},$ (7)
hence we need only to consider the mean magnon number $\left|M\right|^{2}$ as an order parameter to characterize the QPT of the system. In order to derive critical conditions of the phase transition, we first concentrate our attention on $\left|M\right|_{+}^{2}\geq 0$. In the case of $-\tilde{\Delta}_{m}\geq 0$ (i.e., $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta$), we take $\xi^{2}G^{2}-\tilde{\gamma}^{2}=0$ in Eq. (6), which gives rise to $G=G_{c1}$ with
$G_{c1}=\frac{-g^{2}+\sqrt{4\gamma^{2}\tilde{\Delta}_{a}^{2}+\left(g^{2}+2\gamma\kappa\right)^{2}}}{2\gamma}.$ (8)
Thus the critical driving strength $G_{c1}$ is independent of the detuning $\Delta_{m}$. In the other case, when $-\tilde{\Delta}_{m}\leq 0$ and $\sqrt{\xi^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, the critical driving strength is
$G_{c2}=\sqrt{\frac{\left(g^{2}-\tilde{\Delta}_{a}\Delta_{m}\right)^{2}+\gamma^{2}\tilde{\Delta}_{a}^{2}+2g^{2}\gamma\kappa+\gamma^{2}\kappa^{2}+\Delta_{m}^{2}\kappa^{2}}{\gamma^{2}+\Delta_{m}^{2}}}.$ (9)
It clearly see from Eqs. (8) and (9) that both critical driving strengths are subject to the Fizeau shift $\Delta_{F}$. Secondly, we can use the nontrivial solution $\left|M\right|_{-}^{2}$ to obtain the third phase transition point even if it is proved to be unstable in the following. By only considering the case $-\tilde{\Delta}_{m}\geq 0$ and $\sqrt{\xi^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, we obtain the critical driving strength $G=G_{c2}$, which has the same form as Eq. (9).
@@
$\delta\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)\delta a-iJ\delta m-iG\delta a^{\dagger}+\sqrt{2\kappa}\delta a_{in},$
$\delta\dot{m}$ $=$ $-i\left(\tilde{\Delta}_{m}-i\gamma\right)\delta m-iJ\delta a-iKM^{2}\delta m^{\dagger}+\sqrt{2\gamma}\delta m_{in},$
with $\tilde{\Delta}_{m}=\Delta_{m}+2K\left|M\right|^{2}$. Here we have neglected the high-order terms of the fluctuations and $\delta o_{in}$ is the fluctuation on top of $o_{in}$ $\left(o=a,m\right)$. By defining the quadrature operators $\delta Q=\left(\delta a^{\dagger}+\delta a\right)/\sqrt{2}$, $\delta P=i\left(\delta a^{\dagger}-\delta a\right)/\sqrt{2}$, $\delta X=\left(\delta m^{\dagger}+\delta m\right)/\sqrt{2}$, $\delta Y=i\left(\delta m^{\dagger}-\delta m\right)/\sqrt{2}$, Eq. (LABEL:7) can be rewritten in a compact matrix form
$\delta\mathbf{\dot{O}}=\mathbf{U}\cdot\delta\mathbf{O}+\delta\mathbf{O}_{in},$ (11)
with the vector of quadrature components $\delta\mathbf{O=}\left(\delta Q,\delta P,\delta X,\delta Y\right)^{\text{T}}$ and the vector of noise quadratures $\delta\mathbf{O}_{in}=\left(\sqrt{2\kappa}\delta Q_{in},\sqrt{2\kappa}\delta P_{in},\sqrt{2\gamma}\delta X_{in},\sqrt{2\gamma}\delta Y_{in}\right)^{\text{T}}$. The superscript ‘T’ denotes the transpose. The drift matrix $\mathbf{U}$ takes the form

### v2

# Nonreciprocal quantum phase transition in cavity magnonics

@@ Abstract
We investigate the nonreciprocal quantum phase transition in a cavity magnonic system driven by a parametric field, where an yttrium iron garnet (YIG) sphere is placed in a spinning microwave resonator. The system exhibits a rich phase diagram due to both magnon Kerr nonlinearity in YIG and parametric drive on the resonator. Especially, Sagnac-Fizeau shift caused by the spinning of the resonator brings about a significant modification in the critical driving strengths for second- and first-order quantum phase transitions, which means that the highly controllable quantum phase can be realized by the spinning speed of the resonator. More importantly, based on the difference in the detunings of the counterclockwise and clockwise modes induced by spinning direction of the resonator, we show that the phase transition in this system is nonreciprocal, that is, the quantum phase transition occurs when the cavity is driven in one direction but not the other. Our work offers an alternative path to engineer and design nonreciprocal magnonic devices.
@@ I Introduction
In recent years, the cavity magnonic system has become an outstanding platform for exploring light-matter interactions Lachance-Quirion2019 ; Rameshti2022 ; PengYan-2021 . Since ferrimagnetic systems, e.g., yttrium iron garnet (YIG), own high spin density and low damping rate, the strong (even ultra-strong) coupling between magnons in ferrimagnetic systems and photons in microwave cavities can be achieved experimentally Huebl2013 ; Tabuchi2014 ; Zhang2014 ; Goryachev2014 ; Hu-PRL-15 ; Zhang-npjQI-15 . This strong coherent interaction allows one to study many fascinating phenomena, such as magnon dark modes 01 , magnon-photon entanglement 01a , magnon blockade 01b ; 01c ; 01d , non-Hermitian physics 02 ; 01e ; Zhang19 ; 01f ; 01g , quantum states of magnons 01h ; 01i ; 01j ; Zhang23 , cooperative polariton dynamics 01k , and magnon spintronics 01n ; 03 . In addition, it is worth mentioning that the magnon Kerr effect, originating from the magnetocrystalline anisotropy in the YIG, has been theoretically and experimentally demonstrated in cavity-magnon systems 03a ; 03b . The magnon Kerr effect gives rise to the development of some topics, including bistability and tristability of cavity-magnon polaritons 04 ; Nair20 ; Bi21 ; Shen21 ; Bi24 , high-order sideband generation 05 , coherent perfect absorption Zhang-Wang23 , quantum entanglement Yang21 ; Zhang-Agarwal19 , and long-range spin-spin coupling 08 .
Owing to its crucial role in quantum physics and potential applications in quantum technologies, the quantum phase transition (QPT) has attracted considerable attention1 ; 1a ; 2 ; 2a ; 2b ; 3 ; 4 ; 5 ; 6 ; 7 ; 9 ; 10 ; 11 ; 12 ; 13 ; 14 . Different from the classical phase transition happening at a finite temperature, the QPT mainly originates from the quantum fluctuation. In the critical parameter regime, the QPT occurs between two stable phases accompanied by spontaneous symmetry breaking as the temperature tends to the absolute zero of temperature. Recently, QPT has not only been investigated in the Dicke model (or the Tavis-Cummings model) without the ultrastrongly coupling requirement when the cavity field is squeezed via parametric drive 18 , but also been studied in dissipative systems with other types of nonlinearity (e.g., Kerr nonlinearity) 19 ; 20 . With the assistance of magnon Kerr effect and parametric drive, the QPT has also been explored in cavity magnonics 21 ; 22 .
@@
Nonreciprocal physics refers to the phenomenon that a system displays different responses in opposite directions. Up to now, the study of nonreciprocity has been extended to many scientific branches, such as optics 23 ; 24 , acoustics 25 ; 26 ; 27 , and thermodynamics 28 . Traditional nonreciprocal devices break Lorentz reciprocity to achieve nonreciprocity mainly by employing the Faraday effect in magneto-optical crystal materials 29 ; 30 . In such nonreciprocal devices, the used magnetic materials are bulky due to the requirement of high susceptibility to external magnetic field interference, which makes them unsuitable for on-chip integration. To circumvent these obstacles, a variety of nonreciprocity schemes have been proposed based on nonlinear optics 31 ; 32 ; 32a ; 32b , optomechanics 33 ; 34 ; 35 , non-Hermitian optics 36 ; 37 ; 38 ; 39 , etc. In particular, a recent experiment confirmed that an optical diode with 99.6% isolation has been realized by using a spinning resonator 40 , in which the nonreciprocity is caused by the Fizeau shift of circulating lights. Subsequently, there are increasing interests in the connection of the nonreciprocity with other quantum effects, such as quantum blockades 41 ; 42 ; 42a ; 43 ; 44 ; 45 ; 46 ; 47 ; 47a ; 48 , quantum entanglement 49 ; 50 ; 51 ; Chen23 ; 52 , mechanical squeezing 53 ; 54 ; 55 , phonon and magnon laser 56 ; 57 ; 58 , sideband responses 59 ; 60 , single-photon state conversion 60a , and superradiant phase transitions 61c ; 61d ; 61 . However, the nonreciprocal QPT has not yet been investigated in cavity magnonics.
Here we propose a scheme to realize a nonreciprocal QPT in a spinning microwave magnonic system. We first display the phase diagram of the hybrid system, which contains parity-symmetric phase (PSP), parity-symmetry-broken phase (PSBP), and bistable phase (BP). In particular, the Sagnac effect caused by the spinning of the resonator can result in different influences on the detuning of the counterclockwise (CCW) and clockwise (CW) fields compared to the driving field. This leads to the occurrence of the nonreciprocal QPT for driving the resonator from the opposite directions. By investigating the different behaviors of the mean magnon number in the vicinity of the critical point for different Fizeau shifts, we show that the QPT is not only either continuous (second-order) or discontinuous (first-order), but also nonreciprocal for different driving directions. Moreover, an isolation parameter is introduced to quantitatively describe the nonreciprocal QPT. At last the nonreciprocal fluctuation of magnon number is also numerically simulated for displaying the features of the nonreciprocal QPT from the point of view of quantum fluctuation. Nonreciprocal QPT may open up the prospect of engineering nonreciprocal devices for applications in, e.g., on-chip unidirectional quantum sensing 61a and quantum metrology 61b .
In other systems, nonreciprocal phase transition has been investigated 61c ; 61d ; 61 . Especially, Ref. 61c makes a significant contribution to the theory of nonreciprocal phase transition, which originates from asymmetric interactions of multiple species. Subsequently, nonreciprocal coupling is further used to realize nonreciprocal QPT in an open Dicke model 61d . Very recently, Zhu et al 61 proposed to study nonreciprocal QPT in an open dual-coupling Jaynes-Cummings model by utilizing Sagnac effect, where a two-level atom is nonlinearly coupled to modes of a spinning microcavity. As an advantage of this work, it is demonstrated that in the presence of cavity dissipation, the squeezed light can recover the QPT through breaking the requirement of ultrastrong atom-field coupling. On the contrary, although we propose to achieve the nonreciprocity of QPT based on Sagnac effect, our scheme is established in a spinning microwave magnonical system. Here the achievement of QPT relies on not only the squeezed field, but also the magnon Kerr nonlinearity in the YIG. Our scheme exploits a different mechanism to engineer nonreciprocal QPT.
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the theoretical model. Using the Heisenberg-Langevin approach, we give the steady-state solutions of the system. In Sec. III, the nonreciprocal QPT, induced by the Sagnac-Fizeau shift, is studied in detail. Finally, we summarize our results in Sec. IV.
@@ II The model
Figure 1: (a) Schematic of the system. A YIG sphere, biased by a uniform magnetic field $B_{z}$ along the $z$ direction, is placed inside a spinning microwave resonator driven by a squeezed light generated by a flux-driven JPA. (b) Frequencies of the microwave resonator. When the microwave resonator rotates in the fixed counterclockwise (CCW) direction, the resonance frequencies of the CCW mode and clockwise (CW) mode will experience different Fizeau shifts, where the frequencies $\omega_{a}+\left|\Delta_{F}\right|$ and $\omega_{a}-\left|\Delta_{F}\right|$ denote for the CCW and CW modes, respectively.
As schematically shown in Fig. 1, we consider a cavity-magnon system consisting of a spinning resonator and a YIG sphere, where photons in the resonator are coupled to magnons in YIG. In the hybrid system, the resonator is driven by a weak squeezed vacuum field generated by a flux-driven Josephson parametric amplifier (JPA) and the magnetocrystalline anisotropy of YIG results in the Kerr nonlinear interaction among magnons. The Hamiltonian of the total system reads ($\hbar=1$)
@@
where $a$ $\left(a^{\dagger}\right)$ and $m$ $\left(m^{\dagger}\right)$ are, respectively, the annihilation (creation) operators of the cavity and magnon modes with the corresponding resonance frequencies $\omega_{a}$ and $\omega_{m}$. The microwave resonator is parametrically driven with driving amplitude $G$ and pump frequency $\omega_{d}$. $K$ denotes the nonlinear coefficient of the magnon Kerr effect and $J$ is the strength of the cavity-magnon interaction. For a resonator spinning at an angular velocity $\Omega$, the frequencies of the CCW and CW modes experience Sagnac-Fizeau shift, i.e., $\omega_{a}\rightarrow\omega_{a}+\Delta_{F}$, with 40 ; 62
@@
in which $n$ is the refractive index, $r$ is the radius of the resonator, and $\lambda$ $(c)$ is the wavelength (speed) of the light in vacuum. The dispersion term $dn/d\lambda$, characterizing the relativistic origin of the Sagnac effect, is relatively small ($\sim$1%) and thus can be ignored.
In the rotating frame with respect to the driven frequency $\omega_{d}$, the Hamiltonian of the system becomes
$\mathcal{H}$ $=$ $\left(\tilde{\Delta}_{a}-i\kappa\right)a^{\dagger}a+\left(\Delta_{m}-i\gamma\right)m^{\dagger}m+\frac{K}{2}m^{\dagger}m^{\dagger}mm$ (3)
@@
where the dissipations of the cavity-magnon system have been considered. Here $\kappa$ ($\gamma$) is the damping rate of the photon (magnon) mode, $\tilde{\Delta}_{a}=\Delta_{a}+\Delta_{F}$, $\Delta_{a}=\omega_{a}-\omega_{d}/2$, and $\Delta_{m}=\omega_{m}-\omega_{d}/2$. Note that in the absence of YIG sphere (i.e., $J=0$), the squeezing driving can cause the system unstable if $G>\sqrt{\tilde{\Delta}_{a}^{2}+\kappa^{2}}$ 63 . Hence we only consider the case of $G<\sqrt{\tilde{\Delta}_{a}^{2}+\kappa^{2}}$. Interestingly, the non-Hermitian Hamiltonian $\mathcal{H}$ in Eq. (3) commutes with the parity operator $\mathcal{P}=\exp\left[i\pi\left(a^{\dagger}a+m^{\dagger}m\right)\right]$ 2 , i.e., $\left[\mathcal{H},\mathcal{P}\right]=0$, which indicates that the Hamiltonian $\mathcal{H}$ has the parity symmetry. Normally, the total excitation number of the system at steady states can be any natural number. When it takes odd (even) numbers, the steady state of the system has the odd (even) parity. If the total excitation number of the system versus system parameters varies in the even subset $\{0,2,4,6,8,\cdots\}$ or the even subset $\{1,3,5,7,9,\cdots\}$, we say that the corresponding steady state of the system is parity-symmetric. Otherwise, the steady state of the system is parity-symmetry-breaking.
By considering dissipation and noise effects in the Hamiltonian (3) 64 , the quantum Langevin equations (QLEs) describing the dynamics of the hybrid system can be written as
$\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)a-iJm-iGa^{\dagger}+\sqrt{2\kappa}a_{\text{in}},$
$\dot{m}$ $=$ $-i\left(\Delta_{m}-i\gamma\right)m-iKm^{\dagger}mm-iJa+\sqrt{2\gamma}m_{\text{in}},$ (4)
where $a_{\text{in}}$ and $m_{\text{in}}$ are the zero-mean input noise operator for the microwave photon (magnon) mode, respectively. To linearize the Eq. (4), we can expand the operators $a$ and $m$ as a summation of their expectation values and quantum fluctuations, i.e., $a=A+\delta a$ and $m=M+\delta m$. Following from Eq. (4), the dynamical equations for the expectation values $A$ and $M$ are
@@
In the steady-state case ($\dot{A}=\dot{M}=0$), solving Eq. (5) gets three solutions for the mean magnon number $\left|M\right|^{2}$, i.e.,
$\left|M\right|_{0}^{2}=0,\text{ }\left|M\right|_{\pm}^{2}=\frac{-\tilde{\Delta}_{m}\pm\sqrt{\beta^{2}G^{2}-\tilde{\gamma}^{2}}}{K},$ (6)
with $\tilde{\Delta}_{m}=\Delta_{m}-\beta\tilde{\Delta}_{a}$, $\tilde{\gamma}=\gamma+\beta\kappa$, and $\beta=J^{2}/$($\tilde{\Delta}_{a}^{2}+\kappa^{2}-G^{2}$). According to the first equation in Eq. (5), the steady-state photon occupation can be expressed as
$\left|A\right|_{0}^{2}=0,\text{ }\left|A\right|_{\pm}^{2}=\frac{\left(\tilde{\Delta}_{m}+\left|M\right|_{\pm}^{2}\right)^{2}+\gamma^{2}}{g^{2}}\left|M\right|_{\pm}^{2}.$ (7)
Thus, we need only to consider the mean magnon number $\left|M\right|^{2}$ as an order parameter to characterize the QPT of the system. In order to derive critical conditions of the phase transition, we first concentrate our attention on $\left|M\right|_{+}^{2}\geq 0$. In the case of $-\tilde{\Delta}_{m}\geq 0$ (i.e., $\Delta_{m}/\tilde{\Delta}_{a}\leq\beta$), we take $\beta^{2}G^{2}-\tilde{\gamma}^{2}=0$ in Eq. (6), which gives rise to $G=G_{\text{c1}}$ with
$G_{\text{c1}}=\frac{-g^{2}+\sqrt{4\gamma^{2}\tilde{\Delta}_{a}^{2}+\left(g^{2}+2\gamma\kappa\right)^{2}}}{2\gamma}.$ (8)
Thus the critical driving strength $G_{\text{c1}}$ is independent of the detuning $\Delta_{m}$. In the other case, when $-\tilde{\Delta}_{m}\leq 0$ and $\sqrt{\beta^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, the critical driving strength is
$G_{\text{c2}}=\sqrt{\frac{\left(g^{2}-\tilde{\Delta}_{a}\Delta_{m}\right)^{2}+\gamma^{2}\tilde{\Delta}_{a}^{2}+2g^{2}\gamma\kappa+\gamma^{2}\kappa^{2}+\Delta_{m}^{2}\kappa^{2}}{\gamma^{2}+\Delta_{m}^{2}}}.$ (9)
It clearly see from Eqs. (8) and (9) that both critical driving strengths are subject to the Fizeau shift $\Delta_{F}$. Secondly, we can use the nontrivial solution $\left|M\right|_{-}^{2}$ to obtain the third phase transition point even if it is proved to be unstable in the following. By only considering the case $-\tilde{\Delta}_{m}\geq 0$ and $\sqrt{\beta^{2}G^{2}-\tilde{\gamma}^{2}}\geq 0$, we obtain the critical driving strength $G=G_{\text{c2}}$, which has the same form as Eq. (9).
@@
$\delta\dot{a}$ $=$ $-i\left(\tilde{\Delta}_{a}-i\kappa\right)\delta a-iJ\delta m-iG\delta a^{\dagger}+\sqrt{2\kappa}a_{\text{in}},$
$\delta\dot{m}$ $=$ $-i\left(\tilde{\Delta}_{m}-i\gamma\right)\delta m-iJ\delta a-iKM^{2}\delta m^{\dagger}+\sqrt{2\gamma}m_{\text{in}},$
with $\tilde{\Delta}_{m}=\Delta_{m}+2K\left|M\right|^{2}$. Here we have neglected the high-order terms of the fluctuations. By defining the quadrature operators $\delta Q=\left(\delta a^{\dagger}+\delta a\right)/\sqrt{2}$, $\delta P=i\left(\delta a^{\dagger}-\delta a\right)/\sqrt{2}$, $\delta X=\left(\delta m^{\dagger}+\delta m\right)/\sqrt{2}$, and $\delta Y=i\left(\delta m^{\dagger}-\delta m\right)/\sqrt{2}$, Eq. (LABEL:7) can be rewritten in a compact matrix form
$\delta\mathbf{\dot{O}}=\mathbf{U}\cdot\delta\mathbf{O}+\mathbf{O}_{\text{in}},$ (11)
with the vector of quadrature components $\delta\mathbf{O=}\left(\delta Q,\delta P,\delta X,\delta Y\right)^{\text{T}}$ and the vector of noise quadratures $\mathbf{O}_{\text{in}}=\left(\sqrt{2\kappa}Q_{\text{in}},\sqrt{2\kappa}P_{\text{in}},\sqrt{2\gamma}X_{\text{in}},\sqrt{2\gamma}Y_{\text{in}}\right)^{\text{T}}$ with $Q_{\text{in}}=\left(a_{\text{in}}^{\dagger}+a_{\text{in}}\right)/\sqrt{2}$, $P_{\text{in}}=i\left(a_{\text{in}}^{\dagger}-a_{\text{in}}\right)/\sqrt{2}$, $X_{\text{in}}=\left(m_{\text{in}}^{\dagger}+m_{\text{in}}\right)/\sqrt{2}$, and $Y_{\text{in}}=i\left(m_{\text{in}}^{\dagger}-m_{\text{in}}\right)/\sqrt{2}$. The superscript ‘T’ denotes the transpose. The drift matrix $\mathbf{U}$ takes the form